How can i divide the following?
40p9q3
40p8q3

5p8q3
5p9q3

Answer by checkley71(8403)

40P^9Q^3/40P^8Q^3=P/1
-----------------------------
5P^8Q^3/5P^9Q^3=1/P
------------------------------